OpenStudy (anonymous):

I think its substitution.. 3x - y = 1 x + y = 7

OpenStudy (anonymous):

Add the equations together: 3x - y + x + y = 1 + 7 4x = 8 x = 2 Now substitute x = 2 into whichever equation (2 is easier) 2 + y = 7 y = 5 So x = 2, y = 5

OpenStudy (anonymous):

Ohkay that's what i got. ^-^ just wanted to make sure.

OpenStudy (anonymous):

Sure thing. If you want to make sure, just substitute your x and y into both equations and see if they work

OpenStudy (anonymous):

Oh yeah. You may want to close this question ;D